# Simulating the Real World: A Unified Survey of Multimodal Generative Models

### v1

@@ 4 Datasets and Evaluations
In this section, we summarize the commonly-used datasets for 2D, video, 3D, and 4D generation in Table IV. Then, we present a unified and comprehensive summary of evaluation metrics in Table V. For quantitative analysis, we evaluate metrics from two perspectives: 1) Quality: assessing the perceptual quality of the synthesized data, independent of input conditions (e.g., text prompts). 2) Alignment: measuring condition consistency, i.e., how well the generated data matches the user’s intended input. For qualitative analysis, the visual quality of generated results plays a critical role in assessing methods. Therefore, we include some human preference-based indicators as references for more effectively conducting user studies, enabling more convincing qualitative analysis results.
5 Future Directions
@@
Despite significant advancements in 2D, video, and 3D generation, 4D generation still faces substantial challenges. These difficulties are compounded by the complex nature of integrating spatial and temporal dimensions. In 2D generation, key hurdles remain in improving realism and the diversity of generated images. For video generation, modeling long-term temporal dynamics and ensuring smooth transitions between frames are critical challenges. In 3D generation, balancing high-quality output with computational efficiency remains a key issue. Addressing these challenges is essential for advancing 4D generation, as it builds upon these established principles. Below, we outline the primary future directions for 4D generation, illustrating how solving these problems will not only benefit 4D models but also push forward progress in 2D, video, and 3D generation.
Multi-Modal Generation: Generating diverse and plausible 4D content that captures the inherent variability of real-world dynamics is a major challenge. Since real-world scenarios are usually multi-modal, current generative models often struggle to capture this diversity and tend to produce unrealistic results. Techniques like conditional generation or latent space modeling are being explored, but achieving both diversity and realism in multi-modal 4D generation remains an open problem.
@@
Physics and Dynamics Modeling: Realistic 4D generation requires accurate modeling of physical interactions, such as collisions, deformations, and fluid dynamics. Incorporating physics-based constraints into generative models is challenging, as it often involves solving complex differential equations or simulating interactions in real time. Balancing realism with computational efficiency is a open problem.
Generalization Across Scenarios: 4D generation models often struggle to generalize across diverse scenarios, such as different object types, motion patterns, or environmental conditions. This is due to the high variability in dynamic 3D content and the limited diversity in training datasets. Developing models that can adapt to unseen scenarios without extensive retraining remains a challenge.
Control and Editability: Enabling flexible user control over 4D generation, such as specifying motion trajectories or editing dynamic content, is challenging. Current methods often lack fine-grained control, making it difficult to generate content that meets specific requirements. Developing intuitive interfaces and algorithms for interactive 4D editing is an open area of research.
High Computational Cost: 4D generation involves processing and storing vast amounts of data, as it requires modeling both spatial and temporal dimensions. This leads to high computational and memory requirements, making real-time or large-scale 4D generation difficult. Efficient compression techniques and scalable architectures are needed to address these challenges.
6 Conclusions
In this survey, we review recent progress and challenges in simulating the real world across appearance, dynamics, and geometry dimensions through the lens of multimodal generative models. We also summarize the commonly used datasets, their properties, and the corresponding evaluation metrics from various perspectives. Despite significant advancements, challenges in scalability, temporal coherence, and dynamic adaptability still remain. We provided open challenges to guide further researches towards more realistic real-world simulation.

### v2

@@ IV Datasets and Evaluations
In this section, we summarize the commonly-used datasets for 2D, video, 3D, and 4D generation in Table VIII. Then, we present a unified and comprehensive summary of evaluation metrics in Table IX. For quantitative analysis, we evaluate metrics from two perspectives: 1) Quality: assessing the perceptual quality of the synthesized data, independent of input conditions (e.g., text prompts). 2) Alignment: measuring condition consistency, i.e., how well the generated data matches the user’s intended input. For qualitative analysis, the visual quality of generated results plays a critical role in assessing methods. Therefore, we include some human preference-based indicators as references for more effectively conducting user studies, enabling more convincing qualitative analysis results. Besides, we advocate emphasizing the practical challenges in deploying generative models, especially those related to computational efficiency. Many state-of-the-art methods require extensive GPU resources and long inference times, which hinder their accessibility and scalability in real-world applications. These factors are not always reflected in evaluation metrics but are essential when considering deployment in resource-constrained environments or interactive systems. We also encourage future benchmarks to include runtime, memory usage, and training cost to better reflect the practical feasibility of generative models.
V Future Directions
@@
Despite the rapid progress in 2D, video, 3D, and 4D generation, many open questions remain, especially when multiple modalities interact. These difficulties are compounded by the complex nature of integrating spatial and temporal dimensions. In 2D generation, key hurdles remain in improving realism and the diversity of generated images. For video generation, modeling long-term temporal dynamics and ensuring smooth transitions between frames are critical challenges. In 3D generation, balancing high-quality output with computational efficiency remains a key issue. Addressing these challenges is essential for advancing 4D generation, as it builds upon these established principles. Therefore, a clear roadmap is helpful not only for 4D research but also for the related 2D, video, and 3D areas that 4D models build upon.
Multimodal Generation: Generating diverse and realistic 4D content is difficult because real scenes span many modes across space, time, and sensory channels. Findings from multimodal learning highlight three persistent barriers. First, reliable alignment across modalities is required so that the synthesised 4D asset follows the given text, image, or video prompt. Second, well-annotated multimodal datasets are scarce, especially in specialised domains, which limits the coverage of real-world dynamics that current models can learn. Third, existing architectures struggle to scale to the memory and computation cost of modelling high-resolution 3D scenes that evolve over time. Conditional diffusion, compositional latent spaces, and cross-modal contrastive pre-training have improved quality, yet achieving both diversity and fidelity under these constraints remains an open research question.
@@
Efficiency and Controllability: Generating 4D assets entails large spatiotemporal tensors and lengthy optimization cycles that strain GPU memory and compute. These costs increase when score-distillation sampling is used, because its implicit formulation makes the results highly sensitive to control signals and to the diffusion prior. Compact representations and scalable architectures are needed to lower these burdens. Meanwhile, current pipelines still allow only coarse control; precise motion paths and semantic edits remain difficult. Progress will require control-oriented algorithms and clear interactive interfaces.
Fidelity and Diversity: Maintaining both fidelity and diversity in 4D generation remains difficult. For fidelity, shape, texture, and motion must stay consistent across time; otherwise, long sequences suffer from flicker, blurring of fine details, and geometric or photometric drift. Diffusion models used in image-to-4D, video-to-4D, and 3D-to-4D generation pipelines struggle to infer missing views and high-frequency detail from limited input, revealing a need for stronger temporal regularization, view-aware conditioning, and memory-efficient architectures that preserve per-frame quality while ensuring smooth evolution. For diversity, current models generalize poorly across varied objects, motion styles, and environments because dynamic 3D content is highly variable and training datasets remain narrow. Broader data coverage, adaptive conditioning, and modular training strategies will be required to achieve robust generalization, so that high-fidelity synthesis is maintained even in unseen scenarios.
Physics and Dynamics Modeling: Realistic 4D generation must reproduce collisions, elastic and plastic deformation, and fluid flow with physical accuracy, yet embedding such physics into generative models is challenging because it involves solving partial differential equations or simulating interactions in real time. This makes it hard to strike a workable balance between physical fidelity and computational speed. A promising way forward is to pair generative networks with differentiable simulators or neural ordinary differential equation layers, letting gradient‐based training enforce Newtonian laws. Building such coupled systems forms the basis of a world model that encapsulates real-world characteristics, providing a natural benchmark for evaluating physically and dynamically consistent 4D generation. Future research can build on these benchmarks to better simulate realistic interactions while maintaining computational efficiency.
VI Conclusions
In this survey, we have reviewed recent advances in multimodal generative models for simulating the real world, focusing on the intertwined dimensions of appearance, dynamics, and geometry. We categorized existing approaches across 2D, video, 3D, and 4D generation, and discussed their representative methodologies, cross-domain relations, and technical differences, supported by comparative visual examples. Furthermore, we summarized commonly used datasets and evaluation metrics to provide a practical reference for benchmarking.
Despite the rapid progress, fundamental challenges persist, such as the scalability of generative pipelines, temporal consistency in long-range sequences, and adaptability to real-world dynamics. We identified several open research directions, including the need for unified representations across modalities, efficient training with sparse supervision, and the integration of physics-based constraints for enhanced realism.
We hope this survey provides not only a comprehensive overview for newcomers but also a foundation for future research towards more coherent, controllable, and physically grounded multimodal generation systems.